OpenStudy (chris215):

How many excess electrons are there in a typical lightning bolt that has about 16 C of charge? so I used the equation e=1.6(10^-19)(16) and got 2.56×10^(-18) but my answer choices are: 1 x 1020 1 x 1019 10 1 x 1016 what did I do wrong?

OpenStudy (anonymous):

this should help

OpenStudy (chris215):

thanks but idk if A is the right answer do you know how to do this problem?

OpenStudy (anonymous):

no, im sorry but im pretty sure that website tells you

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

he maybe can help

OpenStudy (anonymous):

@paki

OpenStudy (chris215):

ok thanks :)

OpenStudy (anonymous):

http://www.jiskha.com/display.cgi?id=1396904516

OpenStudy (anonymous):

that may help too

OpenStudy (chris215):

It did thatnks so much!

OpenStudy (anonymous):

did you get the answer you were looking for?

OpenStudy (chris215):

yeah it is A

OpenStudy (anonymous):

\[Use,\ Q=n\times e\\ where, Q= total\ charge,\ n= no. of electrons\ , e= charge\ on\ 1\ eletcron\\ you\ wanna\ find, n\\so, n=\frac{ Q }{ e } \\n= \frac{ 16 }{ 1.6\times 10^{-19}} \\1\times \ 10^{20}. Ans\] (excess electrons)